# Towards Robust Audiovisual Segmentation in Complex Environments with Quantization-based Semantic Decomposition

### v2

@@ Abstract
Audiovisual segmentation (AVS) is a challenging task that aims to segment visual objects in videos according to their associated acoustic cues. With multiple sound sources and background disturbances involved, establishing robust correspondences between audio and visual contents poses unique challenges due to (1) complex entanglement across sound sources and (2) frequent changes in the occurrence of distinct sound events. Assuming sound events occur independently, the multi-source semantic space can be represented as the Cartesian product of single-source sub-spaces. We are motivated to decompose the multi-source audio semantics into single-source semantics for more effective interactions with visual content. We propose a semantic decomposition method based on product quantization, where the multi-source semantics can be decomposed and represented by several disentangled and noise-suppressed single-source semantics. Furthermore, we introduce a global-to-local quantization mechanism, which distills knowledge from stable global (clip-level) features into local (frame-level) ones, to handle frequent changes in audio semantics. Extensive experiments demonstrate that our semantically decomposed audio representation significantly improves AVS performance, e.g., +21.2% mIoU on the challenging AVS-Semantic benchmark with ResNet50 backbone. https://github.com/lxa9867/QSD.
@@ 1 Introduction
Prompt-guided video object segmentation (VOS) [31, 18, 63, 61, 69, 68] focuses on the segmentation of objects in videos according to prompts and predict the object classes if required. The use of various modalities of prompts, such as text and audio, has promising implications for multi-modal understanding and user-interactive video applications.
@@ 2 Related Work
Audiovisual segmentation and localization. Audiovisual segmentation (AVS), which was recently introduced [68], aims to segment the objects that produce sound at the time of the image frame. Zhou et al. [68] proposed a method with cross-modal attention to locate the sound source, making it the pioneering work in AVS. Recently, an extended task of AVS, audiovisual semantic segmentation (AVSS), is proposed by Zhou et al. [69] which aims to not only segment the mask of sound sources but also predict the category of each sound source. Due to the semantic entanglement in audio, tackling multi-source AVSS is more challenging than AVS task. Zhou et al. [69] follows the TPAVI module in [68] to conduct audiovisual interaction. More recently, CATR [34] introduces an encoding-decoding framework CATR that presents a novel spatial-temporal audio-video fusion block to fully consider the audio-visual combinatorial dependence in a decoupled and memory-efficient manner. Chen et al. [66] creates a novel benchmark called VPO which augments images from COCO [38] with VGGSound [8] to facilitate the AVS research. Sound source localization (SSL) [45, 46, 52, 26, 50, 9, 2] is a related problem to AVS that aims to locate the regions of sounds in the visual frame. Common SSL methods [5, 4, 17, 52] leverage cross-modal correspondence between audio and visual features to locate sounds, which are then displayed as heatmaps. For instance, Mo et al. [45] leverage multi-level audiovisual contrastive learning to effectively locate the objects. Different from previous methods primarily designed for single-source scenarios, our objective is to address the semantic entanglement present in multi-source audios and explore methods for effective interaction between multi-source audios and videos.
Video Segmentation. Audiovisual segmentation is intricately connected to video object segmentation (VOS) [64, 29, 15, 54, 27, 16, 55, 65] and video semantic segmentation (VSS) [37, 56, 70, 28, 48]. Notably, a most closely related task in the video segmentation domain is referring video object segmentation (R-VOS) [6, 62, 53, 60, 33, 61, 57, 44, 12, 36, 23], which seeks to segment objects in visual frames based on linguistic expressions. In R-VOS, each expression typically refers to a single object, akin to the single-source AVS setting. Recently, Ding et al. [19] introduced the MeViS dataset, expanding the R-VOS task to a more generalized setting that allows referring to multiple objects through a single language expression.
@@ 3.3 Global Decomposition
Method Backbone AVS-Object-Single AVS-Object-Multi
AVS-Sementic
$\mathcal{J}\&\mathcal{F}\uparrow$
$\mathcal{J}\uparrow$
$\mathcal{F}\uparrow$
$\mathcal{J}\&\mathcal{F}\uparrow$
$\mathcal{J}\uparrow$
$\mathcal{F}\uparrow$
@@
LVS [9]
ResNet-18
44.5
37.9
51.9
31.3
29.5
33.0
MSSL [50]
ResNet-18
55.6
44.9
66.3
31.4
26.1
36.3
3DC [42]
3DC
66.5
57.1
75.9
43.6
36.9
50.3
17.3
AOT [65]
ResNet-50
25.4
AVS [69]
ResNet-50
78.8
72.8
84.8
53.6
47.9
57.8
20.2
Bi-Gen [24]
ResNet-50
79.8
74.1
85.4
50.9
50.0
56.8
AVSegFormer [21]
ResNet-50
81.2
76.5
85.9
56.2
49.5
62.8
24.9
${}^{\dagger}$ [34]
ResNet-50
81.0
74.9
87.1
59.4
53.1
65.6
Ours
ResNet-50
81.8 77.6
86.0
61.6 59.6
63.5
46.6
@@
iGAN [43]
Swin-Tiny
69.7
61.6
77.8
48.7
42.9
54.4
SST [20]
SSL
73.2
66.3
80.1
49.9
42.6
57.2
LGVT [67]
Swin-Tiny
81.1
74.9
87.3
50.0
40.7
59.3
AVS [69]
PVT-v2-Base
83.3
78.7
87.9
59.3
54.0
64.5
29.8
Ours
Swin-Tiny
83.9 79.5 88.2 64.0 61.9 66.1 53.4
Table 1: Quantitative comparison to AVS and AVSS methods. normal-$\uparrow$ indicates the larger the better. ${}^{\dagger}$ denotes using additional training data.
@@ 3.5 Loss Function
where $\mathrm{sg}\left[\cdot\right]$ stands for stop-gradient operation. $\mathrm{VQ}(\cdot)$ denotes the vector quantization function, where $\mathrm{VQ}(x)=e_{i}=\arg\min_{e^{k}}\|x-e^{k}\|_{2}\in\mathcal{C}$ and $\mathcal{C}=\{e^{k}\}_{k=1}^{K}$ is the shared codebook. The first term aims to update the codebook. The second and third terms aim to minimize the quantization error by forcing the input vector to be quantized to its closest vector in the codebook.
@@ 4 Experiments
ID Method AVS-Object-Multi
AVS-Sementic
$\mathcal{J}\&\mathcal{F}\uparrow$
$\mathcal{J}\uparrow$
$\mathcal{F}\uparrow$
@@
1 Baseline (w/o decomposition)
[…]
50.1
55.7
33.513.${}_{-13.1}$
2 + TD [10]
[…]
54.5
57.9
38.97.${}_{-7.7}$
3 + Non-Q-based-SD
[…]
56.1
59.1
39.4 ${}_{-6.8}$
4 + Q-based-SD
[…]
57.6
61.8
[…]
5 + Q-based-SD + AVSR
[…]
58.2
61.9
[…]
6 + Q-based-SD + AVSR + LC (ours)
61.6
59.6
63.5
46.6
@@ 4.3 Analysis
Visualization of decomposed semantic space. As shown in Fig. 6, we visualize the semantic space with and without semantic decomposition on the AVS-Semantic dataset using t-SNE [59]. Three types of single-source audios (“man”, “guitar”, “piano”) and two types of multi-source audios (“man+guitar”, “man+piano”) are enrolled. Without decomposition, the multi-source features are highly entangled, presenting less evidence related to single-source semantics. However, after performing semantic decomposition, the “man+guitar” feature presents clear evidence related to its corresponding single-source (“man” and “guitar”) semantics. This is reflected in the proximity of the ”man+guitar” feature to the centroids of its corresponding single-source features. The same applies to the ”man+piano” feature.

### v3

# QDFormer: Towards Robust Audiovisual Segmentation in Complex Environments with Quantization-based Semantic Decomposition

@@ Abstract
Audiovisual segmentation (AVS) is a challenging task that aims to segment visual objects in videos according to their associated acoustic cues. With multiple sound sources and background disturbances involved, establishing robust correspondences between audio and visual contents poses unique challenges due to (1) complex entanglement across sound sources and (2) frequent changes in the occurrence of distinct sound events. Assuming sound events occur independently, the multi-source semantic space can be represented as the Cartesian product of single-source sub-spaces. We are motivated to decompose the multi-source audio semantics into single-source semantics for more effective interactions with visual content. We propose a semantic decomposition method based on product quantization, where the multi-source semantics can be decomposed and represented by several disentangled and noise-suppressed single-source semantics. Furthermore, we introduce a global-to-local quantization mechanism, which distills knowledge from stable global (clip-level) features into local (frame-level) ones, to handle frequent changes in audio semantics. Extensive experiments demonstrate that our semantically decomposed audio representation significantly improves AVS performance, e.g., +21.2% mIoU on the challenging AVS-Semantic benchmark with ResNet50 backbone.
@@ 1 Introduction
Prompt-guided video object segmentation (VOS) [31, 18, 40, 41] focuses on the segmentation of objects in videos according to prompts and predict the object classes if required. The use of various modalities of prompts, such as text and audio, has promising implications for multi-modal understanding and user-interactive video applications.
@@ 2 Related Work
Audiovisual segmentation and localization. Audiovisual segmentation (AVS), which was recently introduced [81], aims to segment the objects that produce sound at the time of the image frame. Zhou et al. [81] proposed a method with cross-modal attention to locate the sound source, making it the pioneering work in AVS. Recently, an extended task of AVS, audiovisual semantic segmentation (AVSS), is proposed by Zhou et al. [82] which aims to not only segment the mask of sound sources but also predict the category of each sound source. Due to the semantic entanglement in audio, tackling multi-source AVSS is more challenging than AVS task. Zhou et al. [82] follows the TPAVI module in [81] to conduct audiovisual interaction. More recently, CATR [34] introduces an encoding-decoding framework CATR that presents a novel spatial-temporal audio-video fusion block to fully consider the audio-visual combinatorial dependence in a decoupled and memory-efficient manner. Chen et al. [71] creates a novel benchmark called VPO which augments images from COCO [43] with VGGSound [8] to facilitate the AVS research. Sound source localization (SSL) [50, 51, 57, 26, 55, 9, 2] is a related problem to AVS that aims to locate the regions of sounds in the visual frame. Common SSL methods [5, 4, 17, 57] leverage cross-modal correspondence between audio and visual features to locate sounds, which are then displayed as heatmaps. For instance, Mo et al. [50] leverage multi-level audiovisual contrastive learning [75, 76, 77, 73, 74, 78] to effectively locate the objects. Different from previous methods primarily designed for single-source scenarios, our objective is to address the semantic entanglement present in multi-source audios and explore methods for effective interaction between multi-source audios and videos.
Video Segmentation. Audiovisual segmentation is intricately connected to video object segmentation (VOS) [69, 29, 15, 59, 27, 16, 60, 70, 36, 38, 37, 79, 80] and video semantic segmentation (VSS) [42, 61, 83, 28, 53]. Notably, a most closely related task in the video segmentation domain is referring video object segmentation (R-VOS) [6, 67, 58, 65, 33, 66, 62, 49, 12, 39, 23], which seeks to segment objects in visual frames based on linguistic expressions. In R-VOS, each expression typically refers to a single object, akin to the single-source AVS setting. Recently, Ding et al. [19] introduced the MeViS dataset, expanding the R-VOS task to a more generalized setting that allows referring to multiple objects through a single language expression.
@@ 3.3 Global Decomposition
Method Backbone AVS-Object-Single AVS-Object-Multi AVS-Sementic
@@
LVS [9] ResNet-18 44.9 37.9 51.9 31.3 29.5 33.0 -
MSSL [55] ResNet-18 55.6 44.9 66.3 31.2 26.1 36.3 -
3DC [47] 3DC 66.5 57.1 75.9 43.6 36.9 50.3 17.3
AOT [70] ResNet-50 - - - - - - 25.4
AVS [82] ResNet-50 78.8 72.8 84.8 52.9 47.9 57.8 20.2
Bi-Gen [24] ResNet-50 79.8 74.1 85.4 53.4 50.0 56.8 -
AVSegFormer [21] ResNet-50 81.2 76.5 85.9 56.2 49.5 62.8 24.9
CATR [34] ResNet-50 81.0 74.9 87.1 59.4 53.1 65.6 -
QDFormer ResNet-50 81.8 77.6 86.0 61.6 59.6 63.5 46.6
@@
iGAN [48] Swin-Tiny 69.7 61.6 77.8 48.7 42.9 54.4 -
SST [20] SSL 73.2 66.3 80.1 49.9 42.6 57.2 -
LGVT [72] Swin-Tiny 81.1 74.9 87.3 50.0 40.7 59.3 -
AVS [82] PVT-v2-Base 83.3 78.7 87.9 59.3 54.0 64.5 29.8
QDFormer Swin-Tiny 83.9 79.5 88.2 64.0 61.9 66.1 53.4
Table 1: Quantitative comparison to AVS and AVSS methods. $\uparrow$ indicates the larger the better. After consulting authors of AVS [82], we reached an agreement to adopt mIoU as a more standardized metric for evaluating AVSS since several previous methods incorrectly compared boundary metric $\mathcal{F}$ with mask metric F-score defined in [82].
@@ 3.5 Loss Function
where $\mathrm{sg}\left[\cdot\right]$ stands for stop-gradient operation. $\mathrm{VQ}(\cdot)$ denotes the vector quantization function, where $\mathrm{VQ}(x)=e_{i}=\arg\min_{e^{k}}\|x-e^{k}\|_{2}\in\mathcal{C}$ and $\mathcal{C}=\{e^{k}\}_{k=1}^{K}$ is the shared codebook. $\lambda_{com}$ is a constant. The first term aims to update the codebook. The second and third terms aim to minimize the quantization error by forcing the input vector to be quantized to its closest vector in the codebook.
@@ 4 Experiments
ID Method AVS-Object-Multi AVS-Sementic
@@
1 Baseline (w/o decomposition) 52.9-8.7 50.1 55.7 33.5-13.1
2 + TD [10] 56.2-5.4 54.5 57.9 38.9-7.7
3 + Non-Q-based-SD 57.6-4.0 56.1 59.1 39.4 -6.8
4 + Q-based-SD 59.7-1.9 57.6 61.8 42.5-4.5
5 + Q-based-SD + AVSR 60.1-1.5 58.2 61.9 44.5-2.1
6 + Q-based-SD + AVSR + LC (ours) 61.6 59.6 63.5 46.6
@@ 4.3 Analysis
Visualization of decomposed semantic space. As shown in Fig. 6, we visualize the semantic space with and without semantic decomposition on the AVS-Semantic dataset using t-SNE [64]. Three types of single-source audios (“man”, “guitar”, “piano”) and two types of multi-source audios (“man+guitar”, “man+piano”) are enrolled. Without decomposition, the multi-source features are highly entangled, presenting less evidence related to single-source semantics. However, after performing semantic decomposition, the “man+guitar” feature presents clear evidence related to its corresponding single-source (“man” and “guitar”) semantics. This is reflected in the proximity of the “man+guitar” feature to the centroids of its corresponding single-source features. The same applies to the “man+piano” feature.